OpenStudy (anonymous):

Find B such that B^3=A B and A are matrices, just tell me what to do?

OpenStudy (anonymous):

wait you are cubing the matrix B?

OpenStudy (anonymous):

finding B , which if you cube it , you will get a known matrix A

OpenStudy (anonymous):

Hmm weird never encountered matrix cubing but anyways this is just matrix multiplication of 3 unknown matrices that are row equivalent

OpenStudy (anonymous):

B*B*B=A

OpenStudy (anonymous):

I would set a generic matrix and set it up with unkowns and carry out matrix multiplication. See if anything familiar pops out or you can distribute something out

OpenStudy (anonymous):

alright, I will try it, thanks

OpenStudy (anonymous):

Well I tried squaring two 2x2 matrix a b c d and I found out that the outcome of matrix multiplication gave me a = d

OpenStudy (ash2326):

@imranmeah91 identity matrix can also be your solution!! \[\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

Ofcfourse lol Identity

OpenStudy (anonymous):

Will that be the only solution? Or is there another way to solve for B for another solution?

OpenStudy (ash2326):

I think with just this info, we can only have identity matrix as a solution.

OpenStudy (anonymous):

nvm